OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and second equations, then the second and third equations. x + y – 3z = –8 2x + 2y + z = 12 3x + y – z = –2 Choices: A.x + y = –8 3x + y = –2 B.7x + 7y = 28 5x + 3y = 10 C.x + y = –8 5x + 3y = 14 D.7x + 7y = 28 3x + y = –2

OpenStudy (anonymous):

to eliminate from the 1st and 2nd you have to multiply the 2nd by 3. Then you have x+y-3z=-8 6x+6y+3z=36 and the result is 7x+7y=28 to eliminate from the 2nd and 3rd you leave it as it is and the result is 5x+3y=10 Therefore the answer is B

OpenStudy (anonymous):

ok thank u :)

OpenStudy (anonymous):

B.7x + 7y = 28 5x + 3y = 10